You are here: Foswiki>Commissioning Web>Com20081111 (13 Nov 2008, AndrewRakich)Edit Attach

AGw Commissioning with IRTC on 10-11 November 2008

Observers: J. Hill (at beginning) and A. Rakich in Tucson
Telescope Operator: W. Wack

Summary

Lost the first 4 hours to a telescope problem.

Details (Times in MST below)

New Stuff

Dave A. reports that an accelerometer mounted on M2 is now being logged. So please note the time of any vibrations that we see. Dave did run the telescope this afternoon.

We are running rolled back irs for LUCIFER compatibility.

Torsten has given us a new version of GCS.

Open

18:15 Open with clear sky. Wind at 6 m/sec. T=-6 degC. D=-8 degC (perhaps covered in ice). SX primary at +2 degC.

18:47 pointto M5_0881 gs=0 ACQUIRE
Nov 10 18:47:06 lbtmu05 irc[1569]: MCSPU error returned - ERROR: Failed to transition between HOLDING<->SLEW_TO_TRACK.

18:50 Started to reset mount encoders, but got interupted by other things.

19:00 Stopall and restart MCSPU (although it was probably fine)

19:05 Telescope won't move from EL=0. It started out as an oil pressure problem. But the more difficult issue seems to be friction in the Elevation Lateral Pads of HBS - similar to the problem last Friday evening.

22:06 Home the AGw motors

22:10 They have the telescope moving again!

22:15 SX Primary is now down to -2.5 degC while reference temperature is down to -4.5 degC and outside ambient is down to -6.5 degC.

22:20 pointto M5_0881 gs=0 ACQUIRE (a few times)

Nov 10 22:20:24 lbtmu05 irc[1899]: MCSPU error returned - ERROR rotator is not movable (Not 'READY' for some reason).

Nov 10 22:21:20 lbtmu05 irc[2042]: MCSPU error returned - ERROR : no tracking polynomials are being received!

Nov 10 22:22:05 lbtmu05 irc[2185]: MCSPU error returned - ERROR : no tracking polynomials are being received!

22:22 Stop and restart PCS.

22:23 pointto M5_0881 gs=0 ACQUIRE

Nov 10 22:24:15 lbtmu05 irc[2331]: left PMC Mirror GoToPositionAbs Command Failed because Chase failed to make tolerance

22:26 cl < home$spiral.cl found at offset 60 60

22:40 IE=-30 CA=-52

22:43 pointto BS9107 gs=0 ACTIVE

1:20 so, apart from anything else I managed to somehow close the log without saving properly, so now summarize the lack of progress tonight as follows:

After pointing to BS9107 found that active optics wouldn't collimate below ~500 nm, ranging up to 1000 nm . Also very funny pupiul images, see /tmp wfsc00019-25 say. Uneven pupil illumination, "fish eggs" etc. all indicative of seeing >3". Took some cubes of IRTC data in /Repository/IRTC_Data/081111 to stack and look for uneven pupil illumination in 30 s exposures.

Had problems with the primary mirror chasing and failing, clear by pressing stop chase and collimate, sometimes several times.

Good news: With a limited work out the new GCS seems to be working OK and giving similar Zernike returns as IDL.

Just as we were about to get the DIMM going for a third measurement of seeing had to close for humidity. Re-opening around 1:05 found that the lateral pad problem had returned and that is currently being worked on.

2:00 am. Maybe back on sky, fingers crossed...

2:01 sending preset to M5_0151 GS 1 in ACQUIRE mode. Worked, though star through cloud and out of focus. Manually adjusted focus 0.1 mm on M2, and adjusted IE CA to -35 -45, then closed for weather.

4:15 still closed for weather

4:45 small break in the weather, going to open

5:01 point to M5_1311 GS 1 in acquire mode to check pointing. Right hand shutter couldn't open, proceeding with one eyelid shut.

5:03 adjust IE CA to -24 -54, Resend preset in ACTIVE mode and collimate in GCS. preset failed couldn't find guide star, going to GS 2.

5:07 acquired guidestar and but too faint for collimating, now pointing's adjusted going to BS9107.

5:15 obscured by clouds, decided to call it a night.

Final IE CA -24 -54

-- JohnHill - 11 Nov 2008
Topic revision: r7 - 13 Nov 2008, AndrewRakich
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback